Consider a bond with a 6% annual coupon and a face value of $1,200 Complete the following table. (Enter your responses rounded to two decimal places) Years to Yeld to Current Maturity Maturity Price 4% 6% 4% 8%
Answer #1

ANSWER:

1) YTM is 4% and n = 3 years

face value = $1,200

annual coupon = 6%

yearly payment = 6% * face value = 6% * 1,200 = $72

pv = yearly payment(p/a,i,n) + face value(p/f,i,n)

pv = 60(p/a,4%,3) + 1,200(p/f,4%,3)

pv = 60 * 2.775 + 1,200 *.889

pv = 166.5 + 1,066.8

pv = 1,233.3

2) YTM is 6% and n = 3 years

face value = $1,200

annual coupon = 6%

yearly payment = 6% * face value = 6% * 1,200 = $72

pv = yearly payment(p/a,i,n) + face value(p/f,i,n)

pv = 60(p/a,6%,3) + 1,200(p/f,6%,3)

pv = 60 * 2.673 + 1,200 *.8396

pv = 160.38 + 1,007.52

pv = 1,167.9

3) YTM is 6% and n = 4 years

face value = $1,200

annual coupon = 6%

yearly payment = 6% * face value = 6% * 1,200 = $72

pv = yearly payment(p/a,i,n) + face value(p/f,i,n)

pv = 60(p/a,6%,4) + 1,200(p/f,6%,4)

pv = 60 * 3.465 + 1,200 *.7921

pv = 207.9 + 950.52

pv = 1,158.42

4) YTM is 4% and n = 6 years

face value = $1,200

annual coupon = 6%

yearly payment = 6% * face value = 6% * 1,200 = $72

pv = yearly payment(p/a,i,n) + face value(p/f,i,n)

pv = 60(p/a,4%,6) + 1,200(p/f,4%,6)

pv = 60 * 5.242 + 1,200 *.7903

pv = 314.52 + 948.36

pv = 1,262.88

5) YTM is 8% and n = 6 years

face value = $1,200

annual coupon = 6%

yearly payment = 6% * face value = 6% * 1,200 = $72

pv = yearly payment(p/a,i,n) + face value(p/f,i,n)

pv = 60(p/a,8%,6) + 1,200(p/f,8%,6)

pv = 60 * 4.623 + 1,200 *.6302

pv = 277.38 + 756.24

pv = 1,033.62